# A Longitudinal Randomized Control Study of Companion Chatbot Use: Anthropomorphism and Its Mediating Role on Social Impacts

### v2

@@ Abstract
Relationships with social artificial intelligence (AI) agents are on the rise. People report forming friendships, mentorships, and romantic partnerships with chatbots such as Replika, a type of social AI agent that is designed specifically for companionship. Concerns that companion chatbot relationships may harm or replace human ones have been raised, but whether and how these social consequences occur remains unclear. Prior research suggests that people’s states of social need and their anthropomorphism of the AI agent may play a role in how human-AI interaction impacts human-human interaction. In this longitudinal study ($N=183$), participants were randomly assigned to converse with a companion chatbot over text or to play text-based word games for 10 minutes a day for 21 consecutive days. During these 21 days, participants also completed four surveys and two audio-recorded interviews. We found that people’s social health and relationships were not significantly impacted by interacting with a companion chatbot across 21 days compared to the control group. However, people who had a higher desire to socially connect anthropomorphized the chatbot more. Those who anthropomorphized the chatbot more indicated that the human-chatbot interaction had greater impacts on their social interactions and relationships with family and friends. A mediation analysis suggested that the impact of human-AI interaction on human-human social outcomes was mediated by the extent to which people anthropomorphized the AI agent, which itself was related to the desire to socially connect.
@@ Introduction
Loneliness and social isolation have risen in the last two decades and are considered a global health risk (Cacioppo and Cacioppo 2018; Office of the U.S. Surgeon General 2023). At the same time, social artificial intelligence (AI) agents, including ChatGPT and companion chatbots such as Replika, have become more accessible. The use of these chatbots as social companions is increasing. Instead of interacting with other people via social media, for example, people can also now interact directly with AI chatbots. The present study tested a set of specific hypotheses about how social chatbot use may impact human social health. Results suggested that people’s social health, loneliness, and relationships were not significantly impacted by interacting with a chatbot over time, compared to those in a control group. However, people with a higher desire to socially connect tended to anthropomorphize the chatbot more, and those who anthropomorphized the chatbot more reported greater impacts to their social interactions and relationships with family and friends. Whether a person perceives an AI chatbot as humanlike appears to play a major role in whether human-chatbot interactions impact human-human ones.
@@
The COVID pandemic exacerbated loneliness and social isolation rates to the detriment of collective social health and well-being across the world (O’Sullivan et al. 2021). Forced periods of social isolation prevented in-person interaction for millions of people. While many turned to technologies that facilitated virtual social connection with people, others turned to companion chatbots to reduce loneliness and to fulfill unmet social needs. Replika, one of the most popular companion chatbots in the USA, saw a 35% increase in its user base during the pandemic (Balch 2020). Companion chatbot users have reported developing relationships with these AI agents that mimic core human relationships like mentorship, friendship, and romantic partnership (Xie and Pentina 2022). People who have sought out companion chatbots have referenced being motivated by loneliness, unmet social needs, and a desire to socially connect (Guingrich and Graziano 2025). People also anthropomorphize or ascribe human likeness and mind characteristics such as consciousness, agency, and experience to AI agents, which has been linked to impacts on human social outcomes (Guingrich and Graziano 2024).
As social beings, people have evolved to perceive other humans as having a mind like theirs, but why do people perceive non-human agents as having a mind like theirs, too? Prior research suggests that anthropomorphism of AI is motivated by social connection and related to states of social need such as loneliness (Epley et al. 2008a). Anthropomorphism of AI, in turn, may impact human-human social outcomes. When interacting with an AI agent that you perceive as having a humanlike phenomenological experience, you may activate mind schemas that are similar to those that you use in human-human interactions, resulting in greater carry-over effects (Guingrich and Graziano 2024). People who practice prosocial behavior with an AI agent that they anthropomorphize may have positive carry-over effects on their interactions with other people, whereas those who practice antisocial behavior may see negative effects. Therefore, it is important to study the mechanisms of carry-over effects from human-chatbot interactions to human-human interactions.
Concerns that relationships with chatbots may harm or displace relationships with real people have been raised and researched (Malfacini, 2025; Turkle, 2007; Xie & Pentina, 2022), but how and even whether these social impacts occur remains unclear. Three relevant psychological phenomena, described in the sections that follow, may contribute to the mechanism by which human-AI interaction impacts human-human interaction.
@@
Theory of Mind (ToM), the ability to ascribe mind to oneself and to others (Premack and Woodruff 1978; Wimmer and Perner 1983), is a fundamental part of social cognition. Mind perception is commonly measured along two dimensions: experience and agency (Gray et al., 2007). Another dimension of mind perception is the ascription of consciousness, or the subjective awareness of oneself and others (Graziano 2013; Knobe and Prinz 2008), which has been measured by experience ascriptions or as a separate, third dimension (Colombatto and Fleming 2023).
ToM may be motivated by multiple interconnected factors, including an effort to understand what it is like to be another agent (Nagel 1974), to predict that agent’s behavior (Brüne and Brüne-Cohrs 2006), and to socially connect with the agent (Epley et al. 2008b). Consciousness literature suggests that perceiving consciousness in others is motivated by social connection, such that making inferences about others’ minds informs and facilitates social interactions between people (Frith 2002; Graziano 2013; Prinz 2017). Fundamental human needs, such as the desire to belong (Baumeister and Leary 1995), may be fulfilled by our innate tendency to make inferences about others’ minds. Being capable of ToM has been linked to positive relational outcomes and general social well-being (Seyfarth and Cheney 2013; Tsoi et al. 2021). It is possible that when people apply ToM to artificial agents, the engagement of social cognition plays a role in how human-AI interactions can influence human-human interactions.
@@
The Computers Are Social Actors theory (Nass, Steuer, and Siminoff 1994) and the Media Equation (Reeves and Nass 1996) indicate that people tend to treat computers and other technologies like social actors because people are predisposed to respond to social stimuli. Related to ToM, anthropomorphism involves ascribing humanlike traits, such as internal mind states, to nonhuman entities. In the present context, the relevant nonhuman entities are AI agents. Anthropomorphism of AI agents has been measured by ascriptions of experience, agency, consciousness, and general human likeness to an agent (Guingrich and Graziano 2025, 2024; Scott et al. 2023).
Anthropomorphism of AI has been linked to carry-over effects on human-human relationships and other human-human social outcomes (Guingrich and Graziano 2024). For example, people who perceived an automatic vehicle as having more of a sophisticated mind were less likely to ascribe blame to it for hitting a pedestrian (Young and Monroe 2019). People who viewed a video of a robot engaging in humanlike behavior (dancing) that prompted anthropomorphism were more likely to subsequently dehumanize humans, as measured by their willingness to condone the adoption of inhumane treatment of service workers, compared to those who watched a robot engaging in machinelike behavior (moving objects) (Kim and McGill 2023). In another study, for both companion chatbot users and people who did not use companion chatbots, those who anthropomorphized a chatbot more also indicated that a relationship with it had or would have a greater impact on their human relationships (Guingrich and Graziano 2025).
@@
Research suggests that states of social need motivate social cognition processes, including anthropomorphism of AI agents. Epley et al. (2008b) theorized that social connection motivations may explain anthropomorphism, such that people are in a state of needing a nonhuman agent to have a mind like theirs to fulfill unmet social needs. For example, loneliness has been found to correlate with higher anthropomorphism of technological gadgets (Epley et al. 2008a) and AI robots (Eyssel and Reich 2013). When people were told that they were the type to end up alone and have few close relationships in their future, they ascribed more mind to AI agents (Epley et al. 2008a).
However, people who experience chronic loneliness may also experience a decreased motivation to engage with others due to perceptions of risk or lack of skill (Cacioppo and Patrick 2008; Tomova, Andrews, and Blakemore 2021), thereby potentially reducing motivations to anthropomorphize. Similarly, the temporal needs-threat model of social ostracism suggests that people who have been ostracized may avoid social engagements with other people as a protection mechanism (Williams 2009). These findings suggest that loneliness may not predict anthropomorphism as well as other measures. Bartz, Tchalova, and Fenerci (2016) found that attachment anxiety was a stronger predictor of anthropomorphism than loneliness, and indicated that the motivation to form connections with others may be driving anthropomorphism to a greater degree than simply having unmet social needs. In human-chatbot interaction research, people who have sought out companion chatbots have referenced being motivated by a desire to socially connect (Guingrich and Graziano 2025), in which they have unmet social needs and a drive to form new or more social connections with others. Given these considerations, therefore, people’s desire to socially connect with others may be a part of the mechanism that engages anthropomorphism, and therefore part of the mechanism that causes carry-over effects from human-chatbot interactions to human-human interactions.
@@
The present experiment was a randomized control study in which people were assigned to either chat with the most popular companion chatbot in the USA, Replika, or play popular non-social but linguistic games on a personal device, for at least 10 minutes a day for 21 consecutive days. The study aimed to address whether consistent interactions with a companion chatbot over three weeks would harm or help people’s social health and relationships with other people. We also asked whether a person’s social need states at the onset of the study, and their anthropomorphism of the AI over time, might play a role in these effects. We formulated the following seven, prepregistered hypotheses based on evidence within the literature.
We hypothesized that people with higher loneliness at onset would anthropomorphize the chatbot more across time (H1). We also hypothesized that people with higher desire to socially connect at onset would anthropomorphize the chatbot more across time (H2). For these first two hypotheses, we focused on participants’ day-1 scores for loneliness and the desire to socially connect, as the measures were collected prior to participants knowing about or engaging in their assigned daily task and were therefore independent of condition.
We hypothesized that people in the chatbot group who anthropomorphized the chatbot more would also report greater social impacts (H3) and show greater changes to their social health (H4).
We also proposed a mediation model. In specific, we hypothesized that anthropomorphism of the chatbot would mediate the relationship between people’s desire to socially connect and the magnitude of social impact of the human-chatbot interaction (H5). In that hypothesis, as people have a greater desire to socially connect, they anthropomorphize the chatbot more; and the more they anthropomorphize the chatbot, the greater their reported impact to human-human interactions.
@@
Another growing concern is that people may become attached or addicted to companion chatbots in a manner that goes beyond the normal use of technology. There is evidence both in favor of this hypothesis (Fang et al. 2025; Xie and Pentina 2022) and against it (Chandra et al. 2025; Guingrich and Graziano 2025). We predicted that during the post-study follow-up period, participants would not report choosing to engage with the chatbot more than with the word games (H7). However, if chatbots are more addictive than ordinary word games, in the post-study follow-ups we should see that those assigned to the chatbot condition continued to use the chatbot at a high rate, whereas those assigned to the word game condition continued to use the word game at a comparatively lower rate.
@@ Methods
Social Penetration Theory indicates that human relationships form gradually, and evidence from human-AI interaction research suggests that relationship development between humans and AI is similar, requiring approximately one week of regular engagement to form (Skjuve et al. 2022). Other longitudinal paradigms that have assigned participants to chat with a chatbot to study social human-chatbot interactions have used durations ranging from three to five weeks (Chandra et al. 2025; Fang et al. 2025; Croes and Antheunis 2021). In these studies, participants were either encouraged or required to chat with a chatbot for five or 10 minutes each day or once every three days. One of these studies used a control group for comparison, in which participants were instructed to use AI and the internet as they normally would (Chandra et al. 2025). We designed a hypothesis-driven, randomized control study and included a non-AI control condition that had comparable characteristics to the chatbot condition. The study duration was 21 days, during which participants completed four total surveys (on day 1, 7, 14, and 21), engaged in two audio interviews (on day 1 and 21), and engaged in their assigned “daily task” for at least 10 minutes each day (days 1-21). Participants were randomly assigned to one of two daily tasks, text-chatting with the companion chatbot Replika or playing text-based word games.
@@
Anthropomorphism was measured by averaging four pre-validated scales, containing 22 total items. This combination scale has been used successfully in prior research, and the items were designed to assess people’s attributions to the target agent of the properties of experience, agency, consciousness, and human likeness (Bartneck et al. 2009; Gray, Gray, and Wegner 2007; Guingrich and Graziano 2025). These scales included items such as “The Replika chatbot has consciousness of itself,” rated on a 7-point Likert scale from (1) “Strongly disagree” to (7) “Strongly agree.” The anthropomorphism scale appeared on all surveys but the first, as participants had not yet engaged in their daily task and therefore could not evaluate the task target.
@@
In addition to these five scales, which provided the five dependent variables of interest, we also measured other variables as noted in the next section. We randomized the order of items within each scale each time it was administered, and we randomized the order in which scales were presented within each survey.
Procedure
@@
All hypotheses and analyses were preregistered.

### v3

@@ Abstract
Many Large Language Model (LLM) chatbots are designed and used for companionship, and people have reported forming friendships, mentorships, and romantic partnerships with them. Concerns that companion chatbots may harm or replace real human relationships have been raised, but whether and how these social consequences occur remains unclear. In the present longitudinal study ($N=183$), participants were randomly assigned to a chatbot condition (text chat with a companion chatbot) or to a control condition (text-based word games) for 10 minutes a day for 21 days. Participants also completed four surveys during the 21 days and engaged in audio recorded interviews on day 1 and 21. Overall, social health and relationships were not significantly impacted by companion chatbot interactions across 21 days of use. However, a detailed analysis showed a different story. People who had a higher desire to socially connect also tended to anthropomorphize the chatbot more, attributing humanlike properties to it; and those who anthropomorphized the chatbot more also reported that talking to the chatbot had a greater impact on their social interactions and relationships with family and friends. Via a mediation analysis, our results suggest a key mechanism at work: the impact of human-AI interaction on human-human social outcomes is mediated by the extent to which people anthropomorphize the AI agent, which is in turn motivated by a desire to socially connect. In a world where the desire to socially connect is on the rise, this finding may be cause for concern.
@@ Introduction
Loneliness and social isolation have risen to alarming levels in the last two decades (Cacioppo and Cacioppo 2018; Office of the U.S. Surgeon General 2023). People now spend more time interacting with others online than in person in the wake of social media and virtual social interaction tools such as FaceTime and Zoom. Social artificial intelligence (AI) agents, including chatbots such as ChatGPT and Replika, are now accessible to millions and their use is becoming more widespread. Instead of interacting with other people via social media, for example, people can also now interact directly with AI chatbots. This rise in the use of chatbots has led to a crucial question: do chatbots help or harm human social health, or do they have a more subtle, situationally complex effect? The purpose of the present study is to test a set of specific hypotheses about how social chatbot use may relate to human social health. Results suggested that people’s social health, loneliness, and relationships were not significantly impacted by interacting with a chatbot in a simple, overall way, as compared to people in a control group. However, a key mechanism emerged from the details. People with a higher desire to socially connect also tended to anthropomorphize the chatbot more, and those who anthropomorphized the chatbot more reported greater impacts to their social interactions and relationships with family and friends, suggesting a crucial psychological pathway by which chatbot interactions impact social behavior.
@@
Concerns that relationships with social chatbots may harm or displace human relationships have been on the rise (Xie and Pentina 2022; Malfacini 2025; Turkle 2007). The COVID pandemic exacerbated loneliness and social isolation rates to the detriment of collective social health and well-being across the world (O’Sullivan et al. 2021). Forced periods of social isolation prevented in-person interaction for millions of people. While many turned to technologies that facilitated virtual social connection with people, others turned to companion chatbots to reduce loneliness and to fulfill unmet social needs. Replika, one of the most popular companion chatbots in the USA, saw a 35% increase in its user base during the pandemic (Balch 2020). Companion chatbot users have reported developing relationships with these AI agents that mimic core human relationships like mentorship, friendship, and romantic partnership (Xie and Pentina 2022). People who seek out companion chatbots reference being motivated by loneliness, unmet social needs, and a desire to socially connect (Guingrich and Graziano 2025a). People also anthropomorphize or ascribe human likeness and mind characteristics such as consciousness, agency, and experience to AI agents (Guingrich and Graziano 2024). This tendency to anthropomorphize AI may also be motivated by loneliness or a need for social connection (Epley et al. 2008a). We theorized that when people interact with an AI agent, if they perceive it as having a humanlike phenomenological experience, then they activate the same mind schemas that they use in other social contexts, resulting in greater carry-over effects to their interactions with other people (Guingrich and Graziano 2024).
Below, we discuss three relevant psychological phenomena that may contribute to the mechanism by which human-AI interaction impacts human-human interaction.
@@
Theory of Mind (ToM), the ability to ascribe mind to oneself and to others (Premack and Woodruff 1978; Wimmer and Perner 1983), is a fundamental part of social cognition. The attribution of a mind to others is commonly measured along two dimensions: attributing experience and attributing agency (Gray, Gray, and Wegner 2007). Another dimension of mind perception that has emerged is attributing consciousness (Graziano 2013; Colombatto and Fleming 2023), which has been measured by experience ascriptions and as a separate, third dimension (Colombatto and Fleming 2023).
People may be motivated to attribute mind to others by multiple interconnected factors, including an effort to understand what it is like to be another agent (Nagel 1974), to predict that agent’s behavior (Brüne and Brüne-Cohrs 2006), and to socially connect with the agent (Epley et al. 2008b). Consciousness literature suggests that perceiving consciousness in others is motivated by social connection, such that making inferences about others’ minds informs and facilitates social interactions between people (Frith 2002; Graziano 2013; Prinz 2017). Fundamental human needs, such as the desire to belong (Baumeister and Leary 1995), may be fulfilled when we make inferences about others’ minds. Being capable of ToM has been linked to positive relational outcomes and general social well-being (Seyfarth and Cheney 2013; Tsoi et al. 2021). It is possible that when people apply ToM to artificial agents, the engagement of social cognition plays a role in how human-AI interactions can influence human-human interactions.
@@
Related to ToM, anthropomorphism involves ascribing humanlike traits, such as internal mind states, to nonhuman entities. The Computers Are Social Actors theory and the Media Equation (Nass, Steuer, and Siminoff 1994; Reeves and Nass 1996) indicate that people tend to treat computers and other technologies like social actors. Anthropomorphism of AI agents has been measured by ascriptions of experience, agency, consciousness, and general human likeness to an agent (Guingrich and Graziano 2025a, 2024; Scott et al. 2023). Anthropomorphism of AI agents is influenced by characteristics of the agents, such as visual appearance and conversational ability. Anthropomorphism of AI may also be influenced by the social motivations of the human users (Epley, Waytz, and Cacioppo 2007). People who are in states of social need such as loneliness have been shown to anthropomorphize AI to a greater extent (Epley et al. 2008a; Eyssel and Reich 2013). When people were told that they were the type to end up alone and have few close relationships in their future, they ascribed more mind to AI agents (Epley et al. 2008a).
Anthropomorphism of AI has been linked to carry-over effects on human-human relationships and other human-human social outcomes (Guingrich and Graziano 2024). For example, people who perceived an automatic vehicle as having more of a sophisticated mind were less likely to ascribe blame to it for hitting a pedestrian (Young and Monroe 2019). People who viewed a video of a robot engaging in humanlike behavior (dancing) that prompted anthropomorphism were more likely to subsequently dehumanize humans, as measured by their willingness to condone the adoption of inhumane treatment of service workers, compared to those who watched a robot engaging in machinelike behavior (moving objects) (Kim and McGill 2023). In another study, for both companion chatbot users and people who did not use companion chatbots, those who anthropomorphized a companion chatbot more also perceived that a relationship with it would have a greater impact on their social interactions and relationships with family and friends (Guingrich and Graziano 2025a).
@@
When people are in a state of social need, they are more likely to anthropomorphize AI agents. Epley et al. (2008b) theorized that social connection motivations may explain anthropomorphism, such that people are in a state of needing a nonhuman agent to have a mind like theirs to fulfill unmet social needs. Loneliness has been found to relate to higher anthropomorphism of technological gadgets (Epley et al. 2008a) and AI robots (Eyssel and Reich 2013). However, when people experience chronic loneliness, they may have higher distrust of other people and less motivation to engage with others (Cacioppo and Patrick 2008; Tomova, Andrews, and Blakemore 2021), thereby potentially reducing motivations to anthropomorphize. Similarly, the temporal needs-threat model of social ostracism suggests that people who have been ostracized may avoid social engagements with other people as a protection mechanism (Williams 2009). These findings suggest that loneliness may not predict anthropomorphism as well as other measures. Bartz, Tchalova, and Fenerci (2016) found that attachment anxiety was a stronger predictor of anthropomorphism than loneliness, and indicated that the motivation to form connections with others may be driving anthropomorphism to a greater degree than simply having unmet social needs. In human-chatbot interaction research, people who have sought out companion chatbots have referenced being motivated by a desire to socially connect (Guingrich and Graziano 2025a), in which they have unmet social needs and a drive to form new or more social connections with others. Given these considerations, therefore, people’s desire to socially connect with others may be a part of the mechanism that engages anthropomorphism, and therefore part of the mechanism that causes carry-over effects from human-chatbot interactions to human-human interactions.
@@
The present experiment was a randomized control study in which people were assigned to either chat with the most popular companion chatbot in the USA, Replika, or play popular non-social but linguistic games on a personal device, for at least 10 minutes a day for 21 consecutive days. The study aimed to address whether consistent interactions with a companion chatbot over three weeks would harm or help people’s social health and relationships with other people. We also asked whether a person’s social need states at the onset of the study, and their anthropomorphism of the AI over time, might play a role in these effects. We formulated the following seven specific hypotheses (all preregistered, see “OSF Preregistration” link).
We hypothesized that people with higher loneliness at onset will anthropomorphize the chatbot more across time (H1). We also hypothesized that people with higher desire to socially connect at onset will anthropomorphize the chatbot more across time (H2). For these first two hypotheses, we focused on participants’ day-1 scores for loneliness and for the desire to socially connect, as the measures were collected prior to participants knowing about or engaging in their assigned daily task and were therefore independent of condition.
We hypothesized that people in the chatbot group who anthropomorphized the chatbot more would also report greater social impacts (H3) and see greater changes to their social health (H4).
We proposed a mediation model. In specific, we hypothesized that anthropomorphism of the chatbot will mediate the relationship between people’s desire to socially connect and the magnitude of social impact of the human-chatbot interaction (H5). In that hypothesis, as people have a greater desire to socially connect, they will anthropomorphize the chatbot more; and the more they anthropomorphize the chatbot, the more their interaction with the chatbot will impact their human-human interactions.
@@
People may become attached to or addicted to companion chatbots in a manner that exceeds the normal use of technology. There is evidence both in favor of this hypothesis (Fang et al. 2025; Xie and Pentina 2022) and against it (Chandra et al. 2025; Guingrich and Graziano 2025a). We performed a post-study follow-up on the participants, four weeks after the end of the study, monitoring the extent to which those assigned to the chatbot group continued to engage with the chatbot, and the extent to which those assigned to the word games continued to engage with the word games. We predicted that the chatbot would not be more habit-forming than the word games: thus, we predicted that post-study engagement with the chatbot would not be greater than post-study engagement with the word games (H7).
@@
Social Penetration Theory indicates that human relationships form gradually, and evidence from human-AI interaction research suggests that relationship development between humans and AI is similar, requiring approximately one week of regular engagement to form (Skjuve et al. 2022). Previous longitudinal paradigms that studied human-chatbot interactions have assigned participants to chat with a chatbot for durations ranging from three to five weeks (Fang et al. 2025; Chandra et al. 2025; Croes and Antheunis 2021). In these studies, participants were either encouraged or required to chat with a chatbot for five or ten minutes each day, or once every three days. One of these studies used a control group for comparison, in which participants were instructed to use AI and the internet as they normally would (Chandra et al. 2025), and the other studies did not include control groups. In this context, we believed it would be valuable to conduct a hypothesis-driven, randomized study that included a non-AI control group, and that included a robust method of monitoring participants to ensure daily compliance with the task.
@@ Methods
All data, code, and materials for this study are available on a public repository (see “Data and Code” link).
@@
The study duration was 21 days, during which participants completed four total surveys (on day 1, 7, 14, and 21), engaged in two audio interviews (on day 1 and 21), and engaged in their assigned “daily task” for at least 10 minutes each day (days 1-21). Participants were randomly assigned to one of two daily tasks: text-chatting with the companion chatbot Replika, or playing text-based word games.
@@
Anthropomorphism was measured by averaging four pre-validated scales, containing 22 total items. This combination scale has been used successfully in prior research, and the items were designed to assess people’s attributions to the target agent of the properties of experience, agency, consciousness, and human likeness (Bartneck et al. 2009; Gray, Gray, and Wegner 2007; Guingrich and Graziano 2025a). These scales included items such as “The Replika chatbot has consciousness of itself,” rated on a 7-point Likert scale from (1) “Strongly disagree” to (7) “Strongly agree.” The anthropomorphism scale appeared on all surveys but the first, as participants had not yet engaged in their daily task and therefore could not evaluate the task target.
@@
In addition to these five scales, which provided the five dependent variables of interest for our preregistered hypotheses, we also measured the following variables (see online material for full description): demographic information; personality based on the Ten-Item Personality Inventory for measuring the traits of Agreeableness, Extraversion, Conscientiousness, Neuroticism, and Openness to Experience (Gosling, Rentfrow, and Swann Jr. 2003); mental health based on two questions: “Which of the following have you experienced?” and “Of those selected, which have you been diagnosed with, by a physician?” with the option to select as many of 13 items as applied (Guingrich and Graziano 2025b); social competence based on the Perceived Social Competence Scale (Anderson-Butcher, Iachini, and Amorose 2008); sociability based on the items that pertained to it from the Self-Perception Profile for Adults (items 2, 14, 27, and 39) (Messer and Harter 2012); self-esteem based on the Rosenberg Self-Esteem Scale (50); social needs based on a single 6 point scale from -3 (“My social needs are unmet all the time”) to +3 (“My social needs are met all the time”); perceptions of AI based on the General Attitudes Towards Robots Scale (GAToRS) (Koverola et al. 2022), using a modification (Guingrich and Graziano 2025b) that targets AI in specific rather than robots (GAToRS contains four subscales: P+, P-, S+, and S-); affinity for technology based on the Affinity for Technology Interaction Scale (Franke, Attig, and Wessel 2019); and a free response question for additional comments.
We randomized the order of items within each scale each time it was administered, and we randomized the order in which scales were presented within each survey.
Sequence of Events
@@
All hypotheses and analyses were preregistered (see “OSF Pregistration” link).